# Improving Hospital Risk Prediction with Knowledge-Augmented Multimodal EHR Modeling Improving Hospital Risk Prediction with Knowledge-Augmented Multimodal EHR Modeling

### v1

@@ 1 Introduction
Appropriate use of clinical prediction tools for early identification of high-risk patients for different conditions allows for clinical decision making, timely interventions, escalation of care, intensive monitoring, and identification of gaps in outpatient management. e.g., (Cai et al. 2016; Kong, Lin, and Hu 2020; Brajer et al. 2020; Kansagara et al. 2011; Mahmoudi et al. 2020). For instance, readmission within a short period is a priority under many regulatory frameworks and value-based care models, where high readmission rates may lead to financial penalties; e.g., (Upadhyay, Stephenson, and Smith 2019; Clement et al. 2017). Therefore, good models of short-term risk prediction (e.g., 30 days) can guide targeted interventions such as more detailed discharge instructions, closer post-discharge monitoring, or referrals to transitional care programs. While traditional prediction tools relied on logistic regression for risk assessment, machine learning methods are increasingly applied to clinical prediction tasks.
@@
There has been a lot of work on developing diverse kinds of machine learning methods for such risk prediction questions using Electronic Health Record (EHR) data, which contain rich information on patient health, e.g., (Cai et al. 2016; Jiang et al. 2025; Qiu et al. 2025; Yang et al. 2023; Brisimi et al. 2018; Bellot and van der Schaar 2019; Cui et al. 2025b; Arsalan 2025; Yu et al. 2024; Wu, Xiao, and Sun 2023; Zhu et al. 2024). Most of this work focuses on structured EHR data, which spans various types and representations (Hartvigsen et al. 2018), including admission/discharge information, procedures and interventions, medications, lab orders and results, billing codes (e.g., ICD, CPT), physiological time-series (e.g., vital signs), imaging data, and clinical documentation such as progress notes or discharge summaries. When used for machine learning approaches to clinical prediction, unstructured data from clinical notes has been used in a fairly simple manner, such as bag of words, TF-IDF (Wu and Zhang 2021; Huang, Osorio, and Sy 2019) to facilitate the use of conventional machine learning methods for clinical tasks. Clinical notes are complex and poorly structured, which limits their utility in clinical informatics tasks, even when using advanced natural language processing (NLP) techniques. While large language models (LLMs) offer a powerful means to process such notes, especially when combined with large biomedical datasets to capture richer semantics beyond keywords and embeddings (Shi et al. 2024; Li et al. 2024; Singhal et al. 2023; Jiang et al. 2025), they still face significant limitations. Hallucinations, factual inaccuracies, and a lack of domain grounding make their use in clinical tasks challenging. Even standard methods such as retrieval-augmented generation (RAG) have produced inaccurate predictions in clinical settings (Shi et al. 2024; Jiang et al. 2025). For instance, models such as Med-PaLM (Singhal et al. 2023) and EHR-Chain-of-Thought (Jiang et al. 2025) exhibit strong language generation capabilities but frequently misinterpret similar-sounding medical terms. Similarly, RAG-based systems like EHRAgent (Shi et al. 2024) struggle when the retrieved context fails to align with a patient’s history. Recent approaches have attempted to enhance LLMs with structured knowledge via graph-based retrieval (e.g., GraphRAG), but their performance remains limited due to a lack of explicit reasoning. Efforts such as MedRetriever (Ye et al. 2021), GraphCare (Jiang et al. 2024), RAM-EHR (Xu et al. 2024), EHR-KnowGen (Niu et al. 2024), and EMERGE (Zhu et al. 2024) inject knowledge into patient representations, yet still fall short of leveraging the full reasoning potential of LLMs. In recent work, Jiang et al. (Jiang et al. 2025) developed KARE, a graphRAG and context augmentation approach, for clinical prediction tasks, which address many challenges associated with using LLMs for clinical tasks on the MIMIC dataset. However, the performance of all prior methods remains limited, and these methods are not usable in clinical decision making.
Challenges. Clinical tasks using EHRs present the following non-trivial challenges: (1) Multi-modal Information: Clinical records in EHRs include both structured and unstructured data. How can we effectively handle both modalities within a single processing pipeline? (2) Long Context: Clinical documents are typically verbose and repetitive. How can we extract relevant information efficiently while avoiding distractions from redundant content? (3) Specialized Medical Domains: EHRs involve domain-specific terminology, such as drugs and treatments. How can we ensure accurate interpretation of such specialized medical information? (4) Highly imbalanced data: Clinical datasets are typically very imbalanced, e.g., about 4% positive for 30-day readmission and about 13% positive for mortality, which has a severe impact on the performance of most ML methods. We will address these questions in the framework Section 4.
Our contributions.
@@
1. We develop a novel framework, KAMELEON (Knowledge-Augmented Multimodal EHR LEarning for Outcome predictioN)
@@
that integrates multimodal EHR data (including structured clinical components and unstructured physician notes), and external biomedical knowledge.
KAMELEON consists of two components (Figure 1):
(1) a graphRAG and knowledge-augmented reasoning approach, which builds on (Jiang et al. 2025), and outputs a prediction and reasoning; and
(2) a structured machine learning module which uses tabular and time-series datasets in the EHR, along with the outputs of the LLM based component.
2. We demonstrate the effectiveness of KAMELEON for two clinical tasks: 30-day readmission risk and mortality prediction, which have been studied extensively, both using MIMIC-III and other private EHRs from specific hospitals.
We compare the performance with a number of structured ML and LLM baselines, with respect to multiple metrics.
KAMELEON achieves significantly better performance for these tasks than all prior work on MIMIC-III datasets.
We observe a relative improvement in the AUROC of approximately for the 30-day clinical readmission prediction task, and around for the mortality prediction task.
Similarly, we achieve a relative AUROC improvement of for 30-day readmission and for in-hospital mortality over the best unstructured LLM baseline (Llama3-Med42-8B). Our work outperforms clinical score methods as well.
The only other prior work which has similar performance for 30-day readmission (Bopche et al. 2024) is on a Norwegian EHR dataset, which is significantly less imbalanced ( readmission positive rate, instead of in the case of MIMIC-III).
@@
In summary, KAMELEON is the first systematic framework to enhance the power of LLMs for healthcare prediction tasks by using graph-guided knowledge retrieval with structured machine learning methods; we expect this framework to be easily applicable to other clinical questions beyond the ones we study here.
Note: Many details of the methodology and additional results are included in the Appendix for brevity.
2 Related Works
Clinical Prediction using Statistical and Machine Learning Methods: Structured data collected from electronic health records (EHRs) contain valuable information about patients and can be instrumental in capturing patient states and predicting medical outcomes (e.g., mortality) (Jiang et al. 2025; Qiu et al. 2025; Yang et al. 2023). To leverage these data, traditional machine learning models—such as linear regression (Brisimi et al. 2018), decision trees (Herazo-Padilla et al. 2021), and Bayesian approaches (Bellot and van der Schaar 2019)—use structured features as input to model the complex relationships between patient features and clinical outcomes. More recently, deep learning–based frameworks have also been explored to capture these complex dynamics and have shown great potential (Cui et al. 2025b; Arsalan 2025; Yu et al. 2024; Wu, Xiao, and Sun 2023; Zhu et al. 2024). However, deep learning-based approaches remain largely restricted to structured data and rely heavily on statistical correlations between input features and outcome labels. Deep learning-based approaches also often lack interpretability grounded in explicit clinical knowledge and tend to be highly specific to the training dataset, limiting their generalizability to other clinical settings (Jiang et al. 2025, 2024).
NLP and Language Models: Unstructured clinical texts (e.g., physician notes, radiology reports, and nursing progress notes) contain rich information about patients (Johnson et al. 2016; Liu et al. 2024; Jiang et al. 2024). Textual documentation includes clinicians’ reasoning and management plans that are not contained elsewhere in the EHR and which may provide direct and high-value insights. To harness this information, various natural language processing (NLP) methods have been proposed for predictive modeling (Mullenbach et al. 2018; Lipton et al. 2016; Yao, Mao, and Luo 2018). Recently, large language models (LLMs) have demonstrated significant progress across a range of tasks, and researchers have begun exploring their potential for processing unstructured clinical text and supporting clinical reasoning (Liu et al. 2024; Niu et al. 2024; Xu et al. 2024). However, one of the major challenges in applying LLMs to healthcare is the risk of errors and hallucinations (Cui et al. 2025a; Shi et al. 2024). To address this, some works incorporate knowledge graphs to explicitly embed clinical knowledge and mitigate LLM inaccuracies (Soman et al. 2024; Jiang et al. 2024). Despite these advances, integrating both structured EHR data and unstructured clinical text in a unified framework remains a challenging problem (Zhang et al. 2021). Existing approaches—such as prompt engineering (Hu et al. 2024), RAG (Zhu et al. 2024), and data fusion (Cui et al. 2024; Thao et al. 2024)—tend to be ad hoc and lack a systematic method for joint modeling.

### v2

@@ 1 Introduction
Appropriate use of clinical prediction tools for early identification of high-risk patients for different conditions allows for clinical decision making, timely interventions, escalation of care, intensive monitoring, and identification of gaps in outpatient management, e.g., [1, 2, 3, 4, 5]. For instance, readmission within a short period is a priority under many regulatory frameworks and value-based care models, where high readmission rates may lead to financial penalties, e.g., [6, 7]. Therefore, effective models of short-term risk prediction (e.g., 30 days) can guide targeted interventions, including more detailed discharge instructions, closer post-discharge monitoring, or referrals to transitional care programs. While traditional prediction tools relied on simple statistical models, e.g., regression and decision trees [8, 9], for risk assessment, more complex machine learning methods are increasingly applied to clinical prediction tasks, e.g., [10, 11, 12, 13, 14]. In this work, we focus on two commonly studied clinical problems: (1) 30-day readmission prediction, which determines whether a patient will be readmitted to the hospital within 30 days after discharge, and (2) mortality prediction, which determines the patient’s in-hospital mortality status.
@@
There has been a lot of work on developing diverse kinds of machine learning methods for these problems using Electronic Health Record (EHR) data, which contain rich information on patient health, e.g., [1, 15, 16, 17, 8, 18, 10, 11, 12, 13, 14]. Most of this work focuses on structured EHR data, which spans various types and representations [19], including admission/discharge information, procedures and interventions, medications, lab orders and results, billing codes (e.g., ICD, CPT), physiological time-series (e.g., vital signs), imaging data, and clinical documentation such as progress notes or discharge summaries. Unstructured data from clinical notes has been used in a fairly simple manner, such as bag of words, TF-IDF [20, 21] to facilitate the use of conventional machine learning methods for clinical tasks. Clinical notes are complex and poorly structured, which limits their utility in clinical informatics tasks, even when using advanced natural language processing (NLP) techniques. While large language models (LLMs) offer a powerful means to process such notes, especially when combined with large biomedical datasets to capture richer semantics beyond keywords and embeddings [22, 23, 24, 15], they still face significant limitations, such as hallucinations, factual inaccuracies, and inadequate domain grounding [22, 15]. For instance, models such as Med-PaLM [24] exhibit strong language generation capabilities but frequently misinterpret similar-sounding medical terms. Recent approaches have attempted to enhance LLMs with structured knowledge via graph-based retrieval (e.g., GraphRAG), but their performance remains limited due to a lack of explicit reasoning, e.g., [25, 26, 27, 28, 14]. In recent work, Jiang et al. [15] developed KARE, a graphRAG and context augmentation approach, for clinical prediction tasks, which address many challenges associated with using LLMs for clinical tasks on the MIMIC dataset.
However, the performance of all prior methods remains limited because clinical tasks using EHRs present many non-trivial challenges: (1) Multi-modality of clinical data: The presence of both structured data and unstructured text requires methods capable of effectively learning from both modalities. (2) Long-context textual data: Clinical text often contains a mix of specialized medical terminology and informal or colloquial expressions, making information retrieval challenging. (3) Severe class imbalance: Prediction tasks are typically highly imbalanced. For example, only about 4% of patients are readmitted within 30 days, resulting in heavily skewed training data.
@@
Here, we develop a novel framework, KAMELEON (Knowledge-Augmented Multimodal EHR LEarning for Outcome predictioN)
@@
that addresses the limitations of prior work by integrating multimodal EHR data (including structured clinical components and unstructured physician notes), and external biomedical knowledge.
KAMELEON consists of two components (as shown in Figure 1):
(1) An unstructured model (Section 4.2.1) that processes clinical notes and retrieves relevant biomedical knowledge using a PubMed-derived graph and knowledge-augmented reasoning, and outputs a prediction for a patient, along with its reasoning; this extends the approach of [15].
Physician notes in EHRs can be lengthy and exceed the context window, and these are summarized using an LLM, and used as context.
To introduce domain-level medical knowledge, we build a biomedical knowledge graph (KG) by combining the Unified Medical Language System (UMLS)[29], PubMed abstracts, and large language model (LLM)-generated entity-relation triples.
KG is partitioned into semantically coherent and well-connected clusters, and textual summary generated by an LLM for the most relevant clusters for each patient cluster, are used to enrich the context.
Furthermore, labeled context is added by identifying semantically similar patient visits, which is used to fine-tune the LLM.
Finally, produces a prediction for the patient, along with a reasoning.
(2) A structured model (Section 4.2.2), which extracts structured features from the patient’s EHR for the stay, including
(a) Static demographic and admission data,
(b) time-varying vitals (which are normalized and summarized, when used as features), and
(c) diagnoses, procedures and medications.
In addition, includes the LLM’s prediction and its tokenized reasoning transformed into an embedding, as inputs.
Finally, different kinds of standard machine learning methods are trained in using these inputs.
We first train separately, and use the LLM outputs to train .
We demonstrate the effectiveness of KAMELEON for the 30-day readmission risk and mortality prediction tasks, which have been studied extensively, both using MIMIC-III and other private EHRs from specific hospitals. We compare the performance with a number of structured ML and LLM baselines, with respect to multiple metrics (discussed in Section 2). KAMELEON consistently outperforms all prior work on MIMIC-III datasets. It also shows clear gains over the strongest unstructured LLM baseline (Llama3-Med42-8B). The only other prior work which has similar performance for 30-day readmission [30] is on a Norwegian EHR dataset, which is significantly less imbalanced ( readmission positive rate, instead of in the case of MIMIC-III).
@@
In summary, KAMELEON is the first systematic framework to enhance the power of LLMs for healthcare prediction tasks through graph-guided knowledge retrieval combined with structured machine learning methods. We expect this framework to be readily applicable to other clinical questions beyond those examined in this study.
| Model | Acc | NPV | Precision (PPV) | Specificity | Sensitivity | Macro F1 | AUROC | AUPRC | ||
| Task: 30-Day Readmission Prediction | ||||||||||
| LogReg (C=1.0) [31] | ✓ | – | 0.831 | 0.869 | 0.036 | 0.951 | 0.013 | 0.464 | 0.463 | 0.090 |
| MLP [30] | ✓ | – | 0.828 | 0.876 | 0.182 | 0.934 | 0.100 | 0.516 | 0.559 | 0.165 |
| LSTM [31] | ✓ | – | 0.820 | 0.876 | 0.163 | 0.925 | 0.100 | 0.512 | 0.569 | 0.152 |
| Claude-3.7-Sonnet† [32] | – | ✓ | 0.240 | 0.790 | 0.199 | 0.068 | 0.927 | 0.227 | 0.498 | 0.199 |
| Medgemma-4b-it† [33] | – | ✓ | 0.350 | 0.770 | 0.190 | 0.270 | 0.690 | 0.350 | 0.480 | 0.190 |
| Llama3-Med42-8B† [34] | – | ✓ | 0.390 | 0.800 | 0.210 | 0.360 | 0.670 | 0.410 | 0.510 | 0.210 |
| – | ✓ | 0.660 | 0.870 | 0.130 | 0.720 | 0.280 | 0.480 | 0.506 | 0.195 | |
| KARE [15] | – | ✓ | 0.271 | 0.785 | 0.191 | 0.131 | 0.851 | 0.269 | 0.491 | 0.191 |
| Task: In-Hospital Mortality Prediction | ||||||||||
| Logistic Regression [31] | ✓ | – | 0.850 | 0.912 | 0.340 | 0.916 | 0.331 | 0.625 | 0.624 | 0.190 |
| LSTM [31] | ✓ | – | 0.690 | 0.800 | 0.260 | 0.800 | 0.250 | 0.530 | 0.560 | 0.240 |
| BRF | ✓ | – | 0.810 | 0.950 | 0.340 | 0.820 | 0.700 | 0.670 | 0.860 | 0.475 |
| LightGBM [30] | ✓ | – | 0.890 | 0.930 | 0.510 | 0.940 | 0.480 | 0.720 | 0.866 | 0.534 |
| MLP [30] | ✓ | – | 0.870 | 0.920 | 0.430 | 0.920 | 0.430 | 0.680 | 0.829 | 0.426 |
| Claude-3.7-Sonnet† [32] | – | ✓ | 0.800 | 0.890 | 0.120 | 0.880 | 0.130 | 0.510 | 0.510 | 0.110 |
| Medgemma-4b-it† [33] | – | ✓ | 0.120 | 0.950 | 0.100 | 0.020 | 0.990 | 0.120 | 0.510 | 0.110 |
| Llama3-Med42-8B† [34] | – | ✓ | 0.160 | 0.950 | 0.120 | 0.100 | 0.970 | 0.190 | 0.530 | 0.120 |
| – | ✓ | 0.614 | 0.890 | 0.134 | 0.641 | 0.413 | 0.474 | 0.527 | 0.125 | |
| KARE [15] | – | ✓ | 0.639 | 0.885 | 0.129 | 0.678 | 0.353 | 0.478 | 0.515 | 0.122 |
| KAMELEON-BRF | ✓ | ✓ | 0.880 | 0.930 | 0.490 | 0.934 | 0.492 | 0.710 | 0.876 | 0.543 |
2 Results
We evaluate our model on the two clinical prediction tasks: in-hospital mortality and 30-day readmission. Our experiments compare performance against strong baselines, including general and medical LLMs, and assess the value of key model components through ablation studies.
2.1 Experimental Setup
Datasets. We use the MIMIC-III dataset [35], which includes structured and unstructured data for over 40,000 ICU patients. It includes structured data (demographics, vitals, labs, admissions, ICD-9-CM codes) and unstructured clinical text (physician notes, discharge summaries, radiology reports). For this study, we focus on physician-authored notes containing clinical reasoning, assessments, and treatment plans. Here, we exclude discharge summaries and notes written after outcomes to prevent label leakage. Only of notes mention hospice, indicating rare explicit terminal indicators. However, a limitation of MIMIC-III for the readmission task is the inability to distinguish planned or elective readmissions and interfacility transfers, which may inflate the count of “avoidable” readmissions.
Biomedical Knowledge. We use abstracts from the annual PubMed Baseline dataset, comprising over 36 million biomedical citation records, to build a medical knowledge graph that enriches LLM input, and reduces hallucinations. We also incorporate UMLS [29] to construct concept-centric subgraphs from EHR data.
Baselines. We compare against Claude 3.5 Sonnet [32], MedGemma [33], LLaMA3-Med [34], and KARE [15], as well as structured-data models including logistic regression, tree-based models, and MLPs used in prior work on MIMIC III [31, 36, 37]. All LLMs are evaluated in a zero-shot setting with the same patient-context prompt. KARE uses similar patient retrieval but lacks clinical notes. Our model incorporates retrieved notes for better context. Implementation details are in the Appendix.
Metrics. We evaluate model performance using a comprehensive set of measures. AUROC and AUPRC capture discrimination ability across thresholds, while overall accuracy reflects the proportion of correct predictions. For class-specific evaluation, we report precision (positive predictive value, PPV), recall (sensitivity), negative predictive value (NPV), and specificity. These capture both the model’s ability to correctly identify positive cases (sensitivity, PPV) and its reliability on negatives (specificity, NPV). Finally, macro F1-score balances precision and recall across classes, ensuring that performance on the minority positive class is not overshadowed by the majority class. All metrics are computed on a held-out test set using standardized preprocessing to ensure comparability across models.
2.2 Performance of KAMELEON
As we will later discuss in Section 4, we explore different kinds of standard ML methods in for making the final prediction, using all the integrated inputs. We refer to the corresponding algorithm as KAMELEON-X, where X represents Balanced-RF (balanced random forest), logistic regression, random forests, LSTM, LightGBM, MLP, or XGBoost.
2.2.1 30-Day Readmission Prediction
As shown in Table 1, our framework with Balanced Random Forest (KAMELEON-BalancedRF) classifier achieves the highest AUROC and notably improves recall on positive cases to , a crucial metric since identifying patients at risk of readmission is clinically imperative. The KAMELEON-MLP model, while achieving the highest overall accuracy and macro F1 , still attains a sensitivity of on positive cases, illustrating the persistent challenge in detecting rare events.
Unstructured LLM-based baselines such as Claude-3.7-Sonnet, MedGemma, LLaMA3-Med, and KARE show substantially lower sensitivity for positives (below 0.3), suggesting that these models struggle to identify the minority class without further fine-tuning or domain-specific adaptation.
In Table 1, KAMELEON-BalancedRF achieves a precision of 0.13, meaning that about one in eight patients flagged as high-risk were actually readmitted. Recall (sensitivity) is 0.79, indicating that the model correctly identifies nearly 8 out of 10 true readmissions—a clinically critical result. The F1-score of 0.55 reflects the balance between precision and recall. The Negative Predictive Value (NPV) is 0.99, showing that almost all patients predicted as low-risk were indeed not readmitted. Specificity is 0.80, meaning the model correctly classifies 8 in 10 patients who were not readmitted as low-risk.
While prior studies like Morgan et al. (2019) [38] reported AUROCs up to 0.81 for readmission, and general models typically ranged from 0.61 to 0.73 [39, 40], our multi-modal approach effectively captures complex clinical nuances.
Relative importance of different classes of inputs.
We conduct an ablation study for the readmission task, where we retrain KAMELEON after dropping different components— the retraining and prediction models from and the demographics used in (Figure 2 and Table 2). We find that the reasoning component output by is very significant, and affects multiple metrics beyond AUROC. In the full model, KAMELEON achieves balanced performance (Acc = 0.80, Macro-F1 = 0.55, AUROC = 0.844, AUPRC = 0.147) with both high specificity (0.80) and sensitivity (0.77). When we drop the reasoning component, sensitivity falls by over 80% (falling to 0.13), and AUPRC is nearly halved, revealing strong bias toward the majority class; however, there is a gain in accuracy (rising to 0.92). Removing the reasoning component from KAMELEON drops performance from 0.844 to 0.7, a 17% decline in AUROC, highlighting the critical role of the finetuned LLM’s reasoning in risk prediction. Removing reasoning and prediction (output from ) causes intermediate degradation, with AUROC falling by , consistent with the loss of calibrated probability signals and semantic rationale. Eliminating reasoning together with demographics and prediction yields the sharpest overall decline, with Macro-F1 dropping by and AUROC by more than (to 0.663), confirming the complementary value of these components. These analyses collectively demonstrate that reasoning substantially improves minority-class detection: it boosts sensitivity and AUPRC by more than threefold compared to variants without it, while also preventing misleading accuracy gains driven solely by the dominant negative class. Figure 2 visualizes these contributions, showing that reasoning, ’s prediction embedding, and demographics each add critical and non-redundant signals for accurate readmission risk estimation.
| Model | Acc | NPV | Precision (PPV) | Specificity | Sensitivity | Macro F1 | AUROC | AUPRC |
| KAMELEON | 0.80 | 0.99 | 0.13 | 0.80 | 0.77 | 0.55 | 0.844 | 0.147 |
| without reasoning | 0.94 | 0.96 | 0.07 | 0.97 | 0.06 | 0.52 | 0.699 | 0.078 |
| without demographic, reasoning, prediction | 0.73 | 0.97 | 0.06 | 0.74 | 0.45 | 0.48 | 0.663 | 0.061 |
| without reasoning, prediction | 0.74 | 0.97 | 0.09 | 0.75 | 0.55 | 0.50 | 0.691 | 0.062 |
Sample Inference Output from .
Table 3 shows a sample output generated by the fine-tuned LLM for the 30-day hospital readmission prediction task. Instead of returning a raw binary value (0 or 1), the model is prompted to generate both a prediction and its reasoning based on the patient’s diagnoses, procedures, medications, and comparisons with similar cases. This structured explanation allows the model to ground its prediction in a clinical context, improving reliability and interpretability over naive classification. By incorporating rationale into the output, the LLM demonstrates better alignment with real-world clinical decision-making.
Patient ID: 25070_0
Patient ID: 93900_1
Analysis of distribution (Readmission).
Figure 3(a) illustrates the distribution of predicted probabilities for the readmission task. In the full model, the majority of negatives cluster near zero, while positives are shifted upward and concentrated above a low threshold of 0.16. This optimal threshold is chosen using Youden’s J statistic, which maximizes sensitivity and specificity. This separation, although not perfectly distinct, reflects the extreme class imbalance in the data: a very low cutoff is required to recover a reasonable fraction of positive cases. The overlap between classes explains the modest AUPRC values, since many positives still lie in regions dominated by negatives. Nevertheless, compared to ablated variants, the full model achieves tighter grouping of positives in the higher-probability region, resulting in better sensitivity and precision–recall trade-offs.
The t-SNE plot (Figure 3(b)) further visualizes patients’ features on the held-out test set. In this space, readmitted patients do not form sharply separated clusters but instead appear partially embedded within the larger manifold of non-readmitted cases. The lack of clear separation reflects the difficulty of the task and the subtlety of signals driving readmission, where positives and negatives overlap substantially. Still, there is evidence of localized groupings of readmitted patients, suggesting that the model captures some latent patterns that distinguish higher-risk subgroups. This partial clustering is consistent with the modest AUPRC values: while the model cannot fully disentangle the classes, it is able to concentrate a portion of true positives in regions of elevated probability. Clinically, this underscores the challenge of predicting readmission but also highlights the value of identifying even partially coherent patient subgroups for targeted follow-up.
2.2.2 In-Hospital Mortality Prediction
Mortality prediction is less imbalanced, with approximately positive cases. Table 1 reports that KAMELEON-XGBoost and KAMELEON-MLP models achieve high accuracy ( and respectively) and AUROC ( and respectively). KAMELEON-XGBoost demonstrates strong performance for mortality prediction, achieving a high precision of , meaning that most patients flagged as high-risk did not survive. It also attains a specificity of and an NPV of , indicating that nearly all patients predicted as low-risk were indeed survivors. Furthermore, for a positive class that, while less imbalanced, is still a minority, the Area Under the Precision-Recall Curve (AUPRC) is a vital metric. Here, the KAMELEON with sets the benchmark, significantly outperforming all other baselines. Unstructured models consistently yield the lowest performance for mortality prediction, AUROC values hover just above random chance (around ), and AUPRCs remain very low (max ), indicating a limited ability to discern between mortality and survival solely from clinical notes. SHAP results (Appendix Figure 4(b)) demonstrate that lab results and vital signs strongly drive predictions, with prediction embeddings playing a less critical role compared to readmission. Overall, Table 1 shows that for both tasks, our multimodal model outperforms all individual structured and unstructured baselines across all metrics.
Relative importance of different classes of inputs.
The SHAP values (Figure 4(b)) show that the prediction embedding has a smaller influence compared to labs and vitals, which contrasts with the readmission task. The AUROC curve (Figure 4(a)) further illustrates that removing reasoning reduces AUROC from 0.92 in the full model to 0.88, a relative drop of about 4%. This indicates that mortality prediction is comparatively easier, as a strong AUROC is retained even without reasoning. The lower class imbalance and the strong signal contained in lab values and diagnostic codes allow the model to capture patterns associated with terminal illness more directly. Nevertheless, reasoning still contributes by improving discrimination at the margin, capturing subtler risk factors that are less apparent in structured features alone.
Sample Inference Output from .
We show inference outputs for two patients on the mortality prediction task, generated by , our finetuned LLM with reasoning in Table 11, Appendix.
Analysis of distribution (In-Hospital Mortality).
Figure 5(a) shows predicted probabilities for in-hospital mortality, separated by true class: survivors (blue) and deceased (red). The dashed line marks the optimal threshold (0.276) balancing sensitivity and specificity.
Class Separability Analysis Using t-SNE (In-hospital Mortality Task).
Figure 5(b) shows a two-dimensional t-SNE embedding of the test data, where the samples are colored by their true class labels (Class 0 in blue and Class 1 in red). The visualization reveals that the majority of the data forms a consistent, structured manifold dominated by Class 0 points, with only a small fraction of Class 1 points distributed across the embedding. Notably, a compact cluster of Class 1 samples appears on the right-hand side, indicating localized patterns that can be exploited by advanced models. This structure suggests that although Class 1 is relatively sparse, it exhibits distinct feature signatures in specific regions, which the proposed KAMELEON‑X model is designed to capture effectively, contributing to improved predictive performance.
2.2.3 Incorporating additional patient context in
Our unstructured model, , extends KARE [15] by incorporating physician notes more explicitly into similar patients, enriching the knowledge graph; in contrast, KARE [15] use only structured EHR data related to drugs, procedures, and conditions as context for similar patients, and the knowledge graph is constructed without considering patient conditions. Our modification ensures that unstructured clinical narratives contribute to prediction alongside structured features, providing a stronger and more comparable baseline.
Our strategy for adding physician notes leads to modest improvements in mortality and more substantial gains in readmission (Table 1). For mortality prediction, AUROC and AUPRC improve by 2.3% and 1.8%, respectively. For readmission, the gains are stronger, with AUROC improving by 3.0% and AUPRC by 2.1%. These results suggest that physician notes provide useful complementary signal, especially for readmission, where unstructured narratives capture behavioral, discharge-related risk factors less visible in structured EHR data.
3 Discussion
In this work, we introduce KAMELEON, a novel framework that effectively integrates multimodal EHR data, including structured clinical features and unstructured physician notes, enhanced by knowledge-augmented LLM reasoning, for robust clinical risk prediction. Our two-stage architecture demonstrated superior performance on both 30-day readmission and in-hospital mortality prediction, with respect to multiple metrics, including the AUROC score. KAMELEON outperforms all prior baselines which only used one type of dataset (structured or unstructured), on most metrics for these two tasks, compared to prior work using the MIMIC-III dataset (Table 1). Multiple types of standard ML methods have been used for these tasks, yet KAMELEON demonstrates clear improvements across all evaluation metrics. None of the currently most powerful LLMs, including a medical LLM trained on clinical data, have comparable performance to KAMELEON. The only exception is that the Llama3-Med42-8B model achieves a higher AUPRC for the readmission task; however, KAMELEON significantly outperforms it across all other metrics.
We find that the reasoning component output by the LLM in , which is used in by constructing an embedding, has high predictive power in both the tasks. For the 30-day readmission task, the embedding constructed using the reasoning output by is very significant— removing this component causes the AUROC to drop from to . This effect is much smaller in the case of the mortality prediction task, but not negligible, dropping the AUROC from 0.92 to 0.88, when this component is dropped. This highlights the synergy achieved by combining these diverse modalities.
This work underscores the significant potential of knowledge-augmented multimodal EHR modeling to enhance early intervention, optimize resource allocation, and improve patient care in complex clinical settings. While LLMs, including medical LLMs trained on specialized data, have a number of limitations in terms of accuracy and hallucinations, their reasoning outputs provide valuable predictive power.
Future work will focus on further validating KAMELEON’s generalizability across diverse clinical settings and exploring its application to a wider range of predictive healthcare tasks. Our framework can be easily extended to other clinical prediction tasks, especially those for which structured models have already been developed. KAMELEON can be applied for such tasks without any changes, and we expect it will provide similar gains.
Scope and Promise for Social Impact. KAMELEON offers a strong opportunity to reduce avoidable hospital readmissions, a major driver of morbidity, cost, and financial penalties [41, 42, 43, 44]. It provides real-time risk predictions for inpatients, enabling more effective discharge planning, case management, and post-acute care. Our model uses real-time patient data to assess readmission risk prior to discharge, supporting individualized case management, discharge planning, census forecasting, and post-acute care coordination. By identifying high-risk patients, our model enables focused use of limited resources, improving efficiency and outcomes.
4 Methods